OpenStudy (anonymous):

1/cuberoot(x-1) = 1/(2/3).... It is algebraically okay to just eliminate the 1/, correct?????

OpenStudy (mathstudent55):

Yes

iYuko (iyuko):

yup

OpenStudy (anonymous):

Ty ty

OpenStudy (mathstudent55):

If \(\dfrac{1}{x} = \dfrac{1}{y}\), then x = y (of course, for x and y not equal to zero.)

OpenStudy (anonymous):

Awesome thanks, my alg so bad lol

OpenStudy (mathstudent55):

You're welcome.